# Synthesizing data products, mathematical models, and observational measurements for lake temperature forecasting

### v1

@@ 2.1 Gaussian process surrogate modeling
Note that $\mathbb{V}\mathrm{ar}\{\boldsymbol{\mu}_{\_}{n_{\_}M}\}<\Sigma_{\_}{n_{\_}M}(\mathcal{X})$, uniformly over all entries of those matrices. [^2] Note also that a non-zero nugget $g$ is present elsewhere in Eq. (5), and when modeling deterministic simulations there is no difference between PIs and CIs. Moreover $\mathbb{V}\mathrm{ar}\{\boldsymbol{\mu}_{\_}{n_{\_}M}\}\rightarrow\boldsymbol{0}$ as $n_{\_}M\rightarrow\infty$ as long as $\mathbf{X}_{\_}{n_{\_}M}$ fills out the input space. In other words, we can learn to emulate $f$ with the surrogate $\hat{f}$ perfectly with enough data. We provide a compare and contrast of PIs and CIs in Figure 6, alongside a description of heteroskedastic modeling in Section 3.2. Finally, it is again worth noting the cubic cost of matrix inversion implied in Eqs. (45) above, even after conditioning on $\hat{\boldsymbol{\theta}}$.
@@ 2.2 Gaussian process surrogates for lake temperatures
To provide an illustration, and to outline how we intend to use GPs for modeling lake temperatures, consider the following setup. Take $f/M$ as GLM and run it for each day of 2021, driven by actual/observed environmental variables in that year, and extract temperatures for ten depths: surface/0 to 9m. So, using the notation established above, $\mathbf{Y}_{\_}{n_{\_}M}$ is comprised of $n_{\_}M=366\times 10=3660$ simulation outputs, and $\mathbf{X}_{\_}{n_{\_}M}$ has $p=2$ columns indexing time $t$ in Julian day, and depth $d$, in meters, respectively. Figure 1 (right) shows a “slice” of some of the elements of this data ($D_{\_}{n_{\_}M}$) as a yellow line where $\mathbf{X}_{\_}{n_{\_}M}[\cdot,2]=6$ meters. (The full $D_{\_}{n_{\_}M}$ is ten times bigger.)
Now consider modeling $D_{\_}{n_{\_}M}$ as a GP, i.e., $\mathcal{GP}(D_{\_}{n_{\_}M})$. Specifically, estimate hyperparameters ($\boldsymbol{\theta}=(\boldsymbol{\gamma},g,\tau^{2})$) by MLE (3), and plug these values $\hat{\boldsymbol{\theta}}$ into predictive equations (4) using $\mathcal{X}=\mathbf{X}_{\_}{n_{\_}M}$, i.e., to evaluate the surrogate at the training locations. Such $n_{\_}M$, in the the small tens of thousands, is on the cusp of manageable with modern architectures and specialized linear algebra libraries such as Intel MKL in about an hour.
Figure 4: The left panel shows the GLM surrogate mean surface $\mu_{\_}{n_{\_}M}(\mathcal{X})$ over day of year (DOY) and depth (m) for 2021. The middle panel shows the mean surface $\boldsymbol{\mu}_{\_}{b}(\mathcal{X})$ of the bias GP $\hat{b}(\cdot)$. The right panel shows a 1D slice of $\boldsymbol{\mu}_{\_}{b}(\mathcal{X})$ where the second column of $\mathcal{X}[,\cdot]=6$ meters, along with 90% PIs corresponding to the slice of $\mathcal{X}[\cdot,2]=6$. For predictions $\hat{f}(\mathcal{X})$ (left) and $\hat{b}(\mathcal{X}$ (middle and right), $\mathcal{X}\equiv\mathbf{X}_{\_}{n_{\_}m}$.
The left panel of Figure 4 shows the resulting surrogate mean surface $\mu_{\_}{n_{\_}M}(\mathcal{X})$ over depth (m) and day of year. This corresponds to the solid blue lines in the right panel of Figure 1 (right), which provide a slice at 6m ( $\mathcal{X}[\cdot,2]=6$) along with 90% PIs corresponding (only) to the slice of $\mathcal{X}[\cdot,2]=6$ meters. Observe in Figure 1 (right) how the surrogate smooths over the GLM simulations and that the intervals have appropriate coverage.
@@
Importantly, $\hat{f}(\mathcal{X})$ in Figures 14 allow us to abstract the raw GLM simulations by smoothing over the noise/jitters that come from noisy or chaotic measurements of environmental variables. Like any statistical model, the surrogate helps us separate signal from noise. Once we have the signal, we can compare it to something else, like actual physical/sensor measurements of lake temperature (black line in Figure 1). From Figure 1, it is obvious that, during the period from April to November, GLM is biased in this application. The surrogate is not essential to make that observation, because the signal and the magnitude of the bias are both high. However, the narrow PIs here provide evidence of a strong bias signal.
The surrogate is more useful in the colder months, when it provides an adequate approximation to observations. Observe that the black lines go outside the blue-dashed PIs in early 2021. This means that, during this period, GLM likely has a warm bias, a signal that would have been hard to separate from noise without the surrogate – in particular without good UQ. Meanwhile, in late 2021, the situation is different: GLM is likely unbiased; any differences with sensor measurements are probably due to noise.
These are qualitative assessments that can be made concrete and quantitative, by using the surrogate to estimate the bias as a function of time and depth. Once estimated, a model for the bias can be used to make a correction (i.e., to obtain a “bias-corrected surrogate”). Let $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ denote the $n_{\_}F$-length vector of sensor observations, or field data, at FCR. In our example, $n_{\_}F=n_{\_}M$ since we have observed temperatures on the same days and at the same depths as our GLM simulations ($\mathbf{X}_{\_}{n_{\_}F}=\mathbf{X}_{\_}{n_{\_}M}$). We could easily have more simulations than sensor observations, say, and we will later in Section 3.
Let $\boldsymbol{\mu}_{\_}{n_{\_}m}\equiv\mu(\mathbf{X}_{\_}{n_{\_}F})$, following Eq. (4) with $\mathcal{X}=\mathbf{X}_{\_}{n_{\_}F}$, which in this case is equivalent to using $\mathcal{X}=\mathbf{X}_{\_}{n_{\_}M}$ (though it need not be). Using this quantity, we may measure the bias of our GLM surrogate relative to the sensor measurements as
@@
We may then form a data set of observed discrepancies between model and field data $D^{b}_{\_}{n_{\_}F}=(\mathbf{X}_{\_}{n_{\_}F},\mathbf{Y}^{b}_{\_}{n_{\_}F})$. Then, regarding bias, or bias correction, $b(\cdot)$ as an unknown function like $f(\cdot)$, we may model it as $\mathcal{GP}(D^{b}_{\_}{n_{\_}F})$ and infer it’s hyperparameters, etc. Readers familiar with the computer modeling literature will recognize a similarity between this setup and the so-called Kennedy and O’Hagan framework (KOH; Kennedy and O’Hagan, 2001) where the two processes: computer model $M$ and field data $F$ are meta-modeled jointly, without directly calculating discrepancies. Our setup here is more modular (Liu et al., 2009), and thus more amenable to up-scaling as described in Section 3. Also, crucially, we are not learning any so-called calibration parameters, although we may explore that in future work (Section 6), so there is no potential for confounding between bias correction and surrogate modeling fits (e.g., Brynjarsdóttir and O’Hagan, 2014).
@@
Predictions obtained via Eq. (4), with moments we shall denote as $\mu_{\_}{n_{\_}F}^{b}(\mathcal{X})$ and $\Sigma_{\_}{n_{\_}F}^{b}(\mathcal{X})$, evaluated at $\mathcal{X}\equiv\mathbf{X}_{\_}{n_{\_}m}$ forming $\hat{b}(\cdot)$ via $\mathcal{GP}(D^{b}_{\_}{n_{\_}F})$, are shown in the middle and right panels of Figure 4. The middle panel shows the full predictive mean surface $\mu_{\_}{n_{\_}F}^{b}(\mathcal{X})$ over depth and DOY, while the right panel shows only the 6m depth slice with 90% PIs. Note that $\mu_{\_}{n_{\_}F}^{b}(\mathcal{X})$ is much more “wiggly” than the surface in the left panel. Focusing on the 6m slice in the right panel of Figure 4, observe the following: in early 2021, the mean predicted bias is small in magnitude and are close to, but do not always contain zero. This indicates a small (and possibly statistically insignificant) warm bias for GLM during this period. Mid-year the magnitude of the predicted bias is large and negative with narrow PIs, indicating substantial warm bias. Finally, later in the year there is a zero/non-significant bias.
@@
Now consider Figure 1 (right), which shows the bias correction, in pink, formed as the sum of two GP fits: GLM surrogate $\hat{f}$ and bias correction $\hat{b}$. One may think of this as (a slice through) the sum of the first two panels in Figure 4, but that would only account for means. A more accurate depiction would involve 2d versions of the slices through the full distribution(s) summarized by the blue lines ($\hat{f}$) in Figure 1 (right) and for the bias correction $\hat{b}$ in the right panel of Figure 4. Mathematically, since $\hat{f}$ and $\hat{b}$ are both MVNs we can follow formulas for sums of MVNs which involve summing the means $(\mu_{\_}{n_{\_}M}(\mathcal{X})+\mu^{b}_{\_}{n_{\_}F}(\mathcal{X})$) and covariances ($\mathbb{V}\mathrm{ar}\{\mu_{\_}{n_{\_}M}(\mathcal{X})\}+\Sigma^{b}_{\_}{n_{\_}F}(\mathcal{X})$) and subtracting off cross-covariances. The pink lines shown in Figure 1 omits subtracted cross-covariances, somewhat overestimating uncertainty.
@@
It is important to note that we are not using $\Sigma_{\_}{n_{\_}M}(\mathcal{X})$, the full predictive variance for $Y(\mathcal{X}\mid D_{\_}{n_{\_}M})$ from Eq. (4) for the bias corrected (co-) variance. Instead we are using the nugget-free one (5) because only the estimated GLM surrogate mean $\boldsymbol{\mu}_{\_}{n_{\_}M}$ was used to define the observed discrepancies, not actual GLM simulations. We shall delve into this further in Section 3.2, and offer a visual in Figure 6 comparing CIs and PIs. For more details, see Gramacy (2020, Sections 5.3.2 & 8.1.3). Observe from Figure 1 (right) that the resulting predictions and PIs in pink are accurate and provide good coverage for the field data. This is perhaps not surprising as we are effectively predicting in-sample, and the GP (or two GPs in fact) provide in a highly flexible non-linear regression. It is their performance out of sample, in particular, as a forecaster for future lake temperatures, which is our focus.
Although we turn to forecasting next in Section 3, it is worth nothing that an analysis such as this is valuable in its own right. Through this exercise, we have gained valuable insight into the nature of the bias in GLM model evaluations that resulted in modifications to the GLM version and configuration. This reduced some of the bias, but not all. In what follows, and especially with our empirical validation in Section 3.3, we utilized that new setup.
@@ 3 GP surrogates for forecasting
There are two challenges to extending the proof-of-concept in Section 2. The first is conceptual: pivoting the modeling framework toward forecasts. We wish to synthesize computer model runs driven by an ensemble of “plausible futures” provided by another computer model, so that they may be compared to what actually happened, as observed in the field. Specifically in the context of lake temperatures, we need a model for how GLM – driven by NOAA ensembles over extended horizons [Figure 3] – relates to temperatures observed by sensors for various depths at FCR.
The second challenge is technological, centering around modeling fidelity and scale. Thirty-day horizons, and 31 NOAA ensemble members forecasting environmental variables, all studied over the course of multiple years (and depths) where observational data has been obtained, represent a very large, and heteroskedastic [Figure 3], simulation campaign. Both, large scale GP modeling (Heaton et al., 2019) and accommodating input-dependent variance for stochastic simulation (Baker et al., 2022), are on the frontier of the current surrogate modeling landscape, as described by recent reviews [cited above]. Those papers outline many choices, most of which are inappropriate in our setting. Here we describe a setup that provides enough flexibility for our NOAA-GLM at FCR while remaining computationally tractable. We are not aware of any examples of where the two (large scale and heteroskedastic GP modeling) have been combined in one application, either generally or in the specific context of forecasting and bias correction. We aim for a practical setup where most of the work is off-loaded to modern libraries so that it is readily extensible to other applications.
@@ 3.1 Modeling a forecasting apparatus
Consider a single horizon of $h$ days into the future, like one day ($h=1$) or one month ($h=30$). We could deploy the setup in Section 2, except with time “shifted” by $t-h$ and with NOAA-simulated environmental variables. More specifically, given NOAA 1–30 day ahead forecasts for each reference date in the training period, GLM will produce 1–30 day ahead lake temperatures at 10 depths. Since the training period is in the past, these faux-future forecasts are what NOAA calls “hindcasts”. [^3] Hindcasts are forecasts for past environmental variables pretending the future (also in the past) is unknown. We may then compare NOAA-GLM hindcasts to what really happened on that day, as recorded in the field data of observed sensor measurements. In other words, we relate the hindcast from time $t-h$ ahead $h$ days to time $t-h+h\equiv t$.
@@
Toward a more holistic model, we propose the following. Let $\mathbf{X}_{\_}{n_{\_}M}$ represent the matrix of inputs as in Section 2, but now augmented with a horizon column indicating $h$, which we take as $h\in\{1,\dots,h_{\_}{\max}=30\}$, an ensemble member column $e\in\{1,\dots,31\}$, and a year column $\lambda$. The $t$ column would be interpreted as the day the forecast was made; the last day for which observational data (i.e., the truth) is known, i.e., $t-h$ from the description above, re-labeled. Columns $e$ and $\lambda$ are merely included to index the data, and we shall provide more detail about how these affect modeling momentarily. We may create such a matrix, i.e., beyond the time and depth columns previously contained in $\mathbf{X}_{\_}{n_{\_}M}$ from Section 2, via Cartesian product: each day and depth pair matched with each horizon, ensemble member and year so that $n_{\_}M\leftarrow n_{\_}M\times 30\times 31\times 3$ compared to the $n_{\_}M$ value in Section 2. Each corresponding entry in the output vector $Y_{\_}{n_{\_}M}$ corresponds to a campaign of NOAA-GLM runs and other configuration inputs.
Continuing to overload notation from Section 2, we have $D_{\_}{n_{\_}M}=(\mathbf{X}_{\_}{n_{\_}M},\mathbf{Y}_{\_}{n_{\_}M})$, relating the time in which a forecast is made, $(t,\lambda)$, to GLM-simulated temperature(s) output at depth $d$ at $h$ horizons into the future, indexed by NOAA ensemble member $e$. Supposing we could surrogate model these simulations, we could again follow the recipe in Section 2 to use their predictive equations to define discrepancies, learn bias corrections, and filter them through to forecasts. There is some nuance to the last few steps, which we outline in Section 3.4, once the surrogate modeling details are in place. However, while it is easy to say “fit a surrogate to $D_{\_}{n_{\_}M}$”, i.e., $\mathcal{GP}(D_{\_}{n_{\_}M})$, it is hard to do. These data are huge, with $n_{\_}M$ in the millions; moreover, NOAA-GLM simulations (across ensemble members) are heteroskedastic [Figure 3]. Addressing these technological challenges are the subject of the next two subsections.

### v2

@@ 2.1 Gaussian process surrogate modeling
Note that $\mathbb{V}\mathrm{ar}\{\boldsymbol{\mu}_{\_}{n_{\_}M}\}<\Sigma_{\_}{n_{\_}M}(\mathcal{X})$, uniformly over all entries of those matrices. [^2] Note also that a non-zero nugget $g$ is present elsewhere in Eq. (5). Moreover $\mathbb{V}\mathrm{ar}\{\boldsymbol{\mu}_{\_}{n_{\_}M}\}\rightarrow\boldsymbol{0}$ as $n_{\_}M\rightarrow\infty$ as long as $\mathbf{X}_{\_}{n_{\_}M}$ fills out the input space. In other words, we can learn to emulate $f$ with the surrogate $\hat{f}$ perfectly with enough data. We compare and contrast of PIs and CIs in Figure 5, alongside our heteroskedastic models in Section 3.2.
@@ 2.2 Gaussian process surrogates for lake temperatures
To illustrate, consider the following. Take $f/M$ as GLM and run it for each day of 2021, driven by observed environmental variables in that year, and extract temperatures for ten depths: surface/0 to 9m. Using the notation established above, $\mathbf{Y}_{\_}{n_{\_}M}$ is comprised of $n_{\_}M=366\times 10=3660$ simulation outputs, and $\mathbf{X}_{\_}{n_{\_}M}$ has $p=2$ columns indexing time $t$ in Julian day, and depth $d$, in meters, respectively. Figure 1 (right) shows a “slice” of some of the elements of this data ($D_{\_}{n_{\_}M}$) as a yellow line where $\mathbf{X}_{\_}{n_{\_}M}[\cdot,2]=6$ meters. (The full $D_{\_}{n_{\_}M}$ is ten times bigger.)
Now consider first modeling $D_{\_}{n_{\_}M}=(\mathbf{X}_{\_}{n_{\_}M},\mathbf{Y}_{\_}{n_{\_}M})$ as a GP, i.e., $\mathcal{GP}(D_{\_}{n_{\_}M})$. Specifically, estimate hyperparameters ($\boldsymbol{\theta}=(\boldsymbol{\gamma},g,\tau^{2})$) by MLE (3), and plug these values $\hat{\boldsymbol{\theta}}$ into predictive equations (4) using $\mathcal{X}=\mathbf{X}_{\_}{n_{\_}M}$, i.e., to evaluate the surrogate at the training locations. Such $n_{\_}M$, in the small tens of thousands, is on the cusp of manageable with modern architectures/libraries.
Figure 4: Left: GLM surrogate mean surface $\mu_{\_}{n_{\_}M}(\mathcal{X})$ over day of year (DOY) and depth (m) for 2021. Middle: mean surface $\boldsymbol{\mu}_{\_}{b}(\mathcal{X})$ of the bias GP $\hat{b}(\cdot)$. Right: 1D slice of $\boldsymbol{\mu}_{\_}{b}(\mathcal{X})$ via f $\mathcal{X}[\cdot,2]=6$ meters, along with 90% PIs.
Figure 4 (left) shows the resulting surrogate mean $\mu_{\_}{n_{\_}M}(\mathcal{X})$ over depth (m) and day of year. This corresponds to the solid blue lines in the right panel of Figure 1, which provides a slice at 6m ($\mathcal{X}[\cdot,2]=6$) along with 90% PIs. Observe in Figure 1 (right) how the surrogate smooths over GLM simulations and that intervals have appropriate coverage.
@@
Importantly, $\hat{f}(\mathcal{X})$ in Figures 14 allow us to abstract raw GLM simulations by smoothing over jitters that come from noisy or chaotic measurements of environmental variables. Like any statistical model, the surrogate helps us separate signal from noise. Once we have the signal, we can compare it to something else, like actual physical/sensor measurements of lake temperature (black line in Figure 1). From Figure 1 we see that GLM is biased from April to November. In fact, narrow PIs here provide evidence of a strong, statistically significant bias since the black lines go outside the blue-dashed PIs in early 2021. This warm bias would have been hard to separate from noise without the surrogate – in particular without good UQ. Meanwhile, in late 2021, the situation is different: GLM is likely unbiased; any differences with sensor measurements are probably due to noise.
These are qualitative assessments that can be made concrete and quantitative by estimating bias as a function of time and depth, and using that estimate to make a correction (i.e., to obtain a “bias-corrected surrogate”). Let $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ denote the $n_{\_}F$-length vector of sensor observations, or field data, at FCR. In our example, $n_{\_}F=n_{\_}M$ since we have observed temperatures on the same days/depths as our GLM simulations ($\mathbf{X}_{\_}{n_{\_}F}=\mathbf{X}_{\_}{n_{\_}M}$). We could easily have more simulations than sensor observations, say, and we will later in Section 3.
Let the predictive mean of $\hat{f}(\mathcal{X})$ be denoted as $\boldsymbol{\mu}_{\_}{n_{\_}M}\equiv\mu_{\_}{n_{\_}M}(\mathcal{X})$, following Eq. (4) with $\mathcal{X}=\mathbf{X}_{\_}{n_{\_}F}$. Using this quantity, we may measure the bias of our GLM surrogate as
@@
We may then form a data set of observed discrepancies between model and field data $D^{b}_{\_}{n_{\_}F}=(\mathbf{X}_{\_}{n_{\_}F},\mathbf{Y}^{b}_{\_}{n_{\_}F})$. Then, by regarding bias, or bias correction, $b(\cdot)$ as an unknown function like $f(\cdot)$, we may model it as $\mathcal{GP}(D^{b}_{\_}{n_{\_}F})$ and infer its hyperparameters, etc. Readers familiar with the computer modeling literature will recognize a similarity between this setup and the so-called Kennedy and O’Hagan framework (KOH; Kennedy and O’Hagan, 2001) where the two processes: computer model $M$ and field data $F$ are meta-modeled jointly, without directly calculating discrepancies. Our setup here is more modular (Liu et al., 2009), and thus more amenable to up-scaling as described in Section 3. Also we are not learning any so-called calibration parameters, although we may explore that in future work (Appendix E), so there is no potential for confounding (e.g., Brynjarsdóttir and O’Hagan, 2014).
@@
Predictions obtained via Eq. (4), with moments $\mu_{\_}{n_{\_}F}^{b}(\mathcal{X})$ and $\Sigma_{\_}{n_{\_}F}^{b}(\mathcal{X})$, evaluated at $\mathcal{X}\equiv\mathbf{X}_{\_}{n_{\_}M}$ forming $\hat{b}(\cdot)$ via $\mathcal{GP}(D^{b}_{\_}{n_{\_}F})$, are shown in the middle and right panels of Figure 4. The middle panel shows full $\mu_{\_}{n_{\_}F}^{b}(\mathcal{X})$ over depth and DOY, while the right panel shows only the 6m depth slice with 90% PIs. Note that $\mu_{\_}{n_{\_}F}^{b}(\mathcal{X})$ is much more “wiggly” than the surface in the left panel. Focusing on the 6m slice, notice that in early 2021 bias is small in magnitude but may be statistically significant. Mid-year bias is large and negative with narrow PIs, indicating substantial warm bias. Finally, later in the year there is zero/non-significant bias.
@@
Now consider Figure 1 (right), which shows the bias correction, in pink, formed as the sum of two GP fits: GLM surrogate $\hat{f}$ and bias correction $\hat{b}$. One may think of this as (a slice through) the sum of the first two panels in Figure 4, but that would only account for means. A more accurate depiction would involve 2d versions of the slices through the full distribution(s) summarized by the blue lines ($\hat{f}$) in Figure 1 (right) and for the bias correction $\hat{b}$ in the right panel of Figure 4. Mathematically, since $\hat{f}$ and $\hat{b}$ are both MVNs we can follow formulas for sums of MVNs which involve summing the means $(\mu_{\_}{n_{\_}M}(\mathcal{X})+\mu^{b}_{\_}{n_{\_}F}(\mathcal{X})$) and covariances ($\mathbb{V}\mathrm{ar}\{\mu_{\_}{n_{\_}M}(\mathcal{X})\}+\Sigma^{b}_{\_}{n_{\_}F}(\mathcal{X})$) and subtracting off cross-covariances. The pink lines shown in Figure 1 omit subtracted cross-covariances, somewhat overestimating uncertainty. The need to subtract cross-covarainces is due to a negative correlation between the GLM surrogate and bias GP (e.g. as one increases, the other would likely decrease) (Kennedy and O’Hagan, 2001; Higdon et al., 2004). However, we cannot access cross-covariances due the R implementation that we use for fitting highly efficient GPs throughout this work (see Section 3.3). Consequently, the total variance of the bias-corrected surrogate is over-inflated slightly. However, we do not believe the over-inflation of this variance negatively affects UQ, as we would prefer conservative UQ over over-confidence of forecasts.
@@
It is important to note that we are not using $\Sigma_{\_}{n_{\_}M}(\mathcal{X})$, the full predictive variance for $Y(\mathcal{X}\mid D_{\_}{n_{\_}M})$ from Eq. (4) for the bias corrected (co-) variance. Instead we are using the nugget-free one (5) because only the estimated GLM surrogate mean $\boldsymbol{\mu}_{\_}{n_{\_}M}$ was used to define the observed discrepancies, not actual GLM simulations. We shall delve into this further in Section 3.2. For more details, see Gramacy (2020, Sections 5.3.2 & 8.1.3). Observe from Figure 1 (right) that the resulting predictions and PIs in pink are accurate and provide good coverage for the field data. This is perhaps not surprising as we are predicting in-sample, and the GP (or two GPs in fact) provide in a highly flexible non-linear regression. It is their performance out of sample, for future lake temperatures, which is our focus.
@@ 3 GP surrogates for forecasting
There are two challenges to extending the proof-of-concept in Section 2. The first is conceptual: pivoting the modeling framework toward forecasts. We wish to synthesize computer model runs driven by an ensemble of “plausible weather futures” so that they may be compared to what actually happened, as observed in the field. We need a model for how GLM – driven by NOAA ensembles over extended horizons [Figure 3] – relates to temperatures observed by sensors at FCR. The second challenge is technological, centering around modeling fidelity and scale. Thirty-day horizons, and 31 NOAA ensemble members over the course of multiple years (and depths) where observational data has been obtained, represent a very large, and heteroskedastic [Figure 3], simulation campaign. Both, large scale GP modeling (Heaton et al., 2019) and accommodating input-dependent variance for stochastic simulation (Baker et al., 2022), are on the frontier of the surrogate modeling landscape. Here we describe a setup that provides enough flexibility for our NOAA-GLM at FCR while remaining practical, off-loading work to modern libraries as much as possible, and computationally tractable.
@@ 3.1 Modeling a forecasting apparatus
Consider a single horizon of $h$ days into the future, like one day ($h=1$) or one month ($h=30$). We could deploy the setup in Section 2, except with time “shifted” by $t-h$ and with NOAA-simulated environmental variables. More specifically, given NOAA 1–30 day ahead forecasts for each reference date in the training period, GLM will produce 1–30 day ahead lake temperatures at 10 depths. Since the training period is in the past, these faux-future forecasts are what NOAA calls “hindcasts”. [^3] Hindcasts are forecasts for past environmental variables pretending the future (also in the past) is unknown. We may then compare NOAA-GLM hindcasts to what really happened on that day; i.e., we relate the hindcast from time $t-h$ ahead $h$ days to sensor measurements at time $t-h+h\equiv t$.
@@
Toward a more holistic model, we propose the following. Let $\mathbf{X}_{\_}{n_{\_}M}$ represent the matrix of inputs as in Section 2, but now augmented with a horizon column $h\in\{1,\dots,h_{\_}{\max}=30\}$, an ensemble member column $\xi\in\{1,\dots,31\}$, and a year column $\lambda$. The $t$ column would be interpreted as the day the forecast was made; the last day for which observational data (i.e., the truth) is known, i.e., $t-h$ from the description above, re-labeled. Columns $\xi$ and $\lambda$ are merely included to index the data, and we shall provide more detail about how these affect modeling momentarily. We may create such a matrix, i.e., beyond the time and depth columns previously contained in $\mathbf{X}_{\_}{n_{\_}M}$ from Section 2, via Cartesian product so that $n_{\_}M\leftarrow n_{\_}M\times 30\times 31\times 3$. Now let $D_{\_}{n_{\_}M}=(\mathbf{X}_{\_}{n_{\_}M},\mathbf{Y}_{\_}{n_{\_}M})$ collect simulation data relating time, horizon, depth, and ensemble member to GLM-simulated temperature(s). Supposing we could surrogate model these simulations, we could again follow the recipe in Section 2, filtering surrogate fits and bias corrections through to forecasts. But there are logistical challenges with $n_{\_}M$ in the millions and with heteroskedastic NOAA-GLM runs [Figure 3].